# Goofy transformations and the hierarchy problem

### v1

@@ Abstract
All of these solutions require a breaking of the protecting symmetry in order to be phenomenologically viable in the sense that they need to allow for electroweak symmetry breaking and the observed Higgs mass. The mechanism of breaking typically renders the Higgs mass computable and introduces deviations in Higgs phenomenology as compared to the plain SM. Also in goofy situations phenomenological consequences are computable and consequences of specific models should lead to testable predictions at future colliders and Higgs factories.
@@
Seemingly as a model building choice, one possibility to compensate the unusual transformation behavior of gauge-kinetic terms under goofy transformations is by introducing an equally goofy transformation of space-time and gauge fields [1]. The combined transformation then acts as a regular symmetry, but only leaves invariant the full effective potential if one simultaneously performs a sign flip of the squared UV cutoff, or squared $\overline{\mathrm{MS}}$ renormalization scale [7, 4] (see also [8, 9]). This might be important to take into account for the “intrinsic” hierarchy problem of the SM [10], as a spurious transformation behavior of the cutoff under goofy transformations can determine whether or not, and at what order, it contributes to the otherwise prohibited Higgs quadratic term.
An alternative route to understand the all-order RG stability of the goofy parameter relations, independent of exotic space-time transformations, stems from the observation that the breaking of goofy symmetry by the kinetic terms can be soft, in the sense that it does not enter quantum corrections of other goofy non-invariant operators, which also explains their RG stability to all orders [3]. In this picture, nothing has to be imposed about the transformation behavior of the RG scale or cutoff, and physical scales can exist which do not have the correct spurion transformation behavior to enter the Higgs quadratic term, hence, giving way for a stabilization of a hierarchy of scales.
@@
$\left(\begin{matrix}H\\ \widetilde{H}\end{matrix}\right)\mapsto\left(\begin{matrix}0&\mathrm{e}^{\mathrm{i}\beta}\\ -\mathrm{e}^{-\mathrm{i}\beta}&0\end{matrix}\right)\left(\begin{matrix}H\\ \widetilde{H}\end{matrix}\right)\,,$ (2)
for flavor- or CP-type transformations, respectively  [^2] It is left implicit throughout this work, but for CP-type transformation we always additionally perform the transformation $(t,x)\mapsto(t,-x)$ as well as the usual action of the complex conjugation outer automorphism on gauge fields and fermions [25, 26]. For fermions the CP transformation on flavor and spinor indices reads, for example, $Q_{\mathrm{L}}\mapsto C_{Q_{\mathrm{L}}}\,\mathcal{C}\,Q_{\mathrm{L}}^{*}$, where $\mathcal{C}$ is the charge conjugation matrix and $C_{Q_{\mathrm{L}}}$ a possible flavor-space transformation.. The phases $\alpha$ and $\beta$ here can always be modified by performing an additional global hypercharge transformation and we often adopt a convention such as to set $\alpha=\beta=0$ for the Higgs transformation  [^3] We remark that an independent passive field redefinition of $H$ and $\widetilde{H}$ cannot be used to absorb the phases in Eq. (1) since such an operation commutes with this transformation. Passive field redefinitions, however, can be used to absorb the phases of the CP-type transformation in Eq. (2), which can be used to show that it is basis-change equivalent to an order 222 transformation. Moreover, passive “goofy” field redefinitions can be used to change to a basis with different sign for the kinetic term..
@@
Extending goofy transformations to fermions, there is a qualification necessary, depending on if one requests presence of additional space-time symmetries or not. For the scalar kinetic term, the requirement of hermiticity in the case of no additional space-time transformation imposes the same requirement as demanding invariance in the presence of additional space-time transformations $\partial_{\mu},\,A^{a}_{\mu}\mapsto-\mathrm{i}\partial_{\mu},\,-\mathrm{i}A^{a}_{\mu}$, as suggested by [1]. For the fermion kinetic term, the two requirements do not align. For example, consider a generic fermion kinetic term (eventually we will take $\Psi=Q_{\mathrm{L}},\,u_{\mathrm{R}},\,d_{\mathrm{R}}$)
@@
where we have included $\times 3$ 3 unitary transformation matrices in flavor space $A_{\Psi}$, $B_{\Psi}$, $C_{\Psi}$, and $D_{\Psi}$  [^6] See [3] for details in the analogous case of scalars and keep in mind footnote [12]..
Starting from the canonical basis, and without additional exotic space-time transformations, hermiticity of the kinetic term requires $B_{\Psi}^{\dagger}A_{\Psi}=\pm\mathbbm{1}$ or $D_{\Psi}^{\dagger}C_{\Psi}=\pm\mathbbm{1}$ for flavor universal regular (+)(+)( + ) or goofy ()(-)( - ) transformations, respectively. Different transformations for different $\Psi$’s, and also flavor non-universal transformations are possible. By contrast, superimposing an additional exotic space-time symmetry $\partial_{\mu},A_{\mu}\mapsto-\mathrm{i}\partial_{\mu},-\mathrm{i}A_{\mu}$, the overall invariance of the kinetic term requires $B_{\Psi}^{\dagger}A_{\Psi}=\mathrm{i}\mathbbm{1}$ or $D_{\Psi}^{\dagger}C_{\Psi}=\mathrm{i}\mathbbm{1}$, which in this case has to hold universally for all $\Psi=Q_{\mathrm{L}},\,u_{\mathrm{R}},\,d_{\mathrm{R}}$. Obviously, these are very different constraints that can lead to different possibilities regarding the possibly allowed transformations and the resulting phenomenology.
For the case of no extra space-time symmetry there is a lot of model building freedom. Not all fermion kinetic terms have to experience a simultaneous sign flip. Which fermion kinetic terms explicitly break the goofy transformations depends on the precise charge assignments under the transformation.
@@
can ensure invariant Yukawa couplings, while the Higgs mass term is prohibited. The transformation is explicitly broken by the $H$, $u_{\mathrm{R}}$ and $d_{\mathrm{R}}$ kinetic terms, which is not automatically bad since such a breaking can be soft, and can control radiative corrections potentially in phenomenologically interesting ways (see [3]).
@@
Let us also consider the case of imposing the additional space time symmetry. If this is combined with the transformations (1) or (2) and demanding invariance of the fermion kinetic terms, the flavor universal fermion goofy transformations are limited to
@@
Discussion.– All our examples show how goofy transformations can be invoked to prohibit the Higgs quadratic term while permitting SM Yukawa couplings. Even if a goofy transformation is explicitly broken by some kinetic terms, it can give rise to all-order stable RG stable fixed points, including mass terms. This has been explicitly shown for some cases of the 2HDM in [1, 3], where RG stable fixed points exist even if the canonical kinetic terms explicitly (softly) break the corresponding goofy transformation  [^8] We add to [3] that for all cases with exact protection of mass terms, one can always find a full rank flavor non-universal (hence non-canonical) kinetic term which is actually conserved under a (in that case Higgs flavor dependent) goofy transformation. Needless to say that it would also be extremely interesting to explore such flavor dependence in fermion kinetic terms.. Irrespectively of whether or not a goofy transformation is actually imposed, the transformation acts as an outer automorphism and, therefore, provides a sufficient condition for a symmetry of the system of beta functions that can be larger than the symmetries of the effective action [3, 6].
@@
This implies that a vanishing beta function of $\mu_{H}$ can be stabilized to all order by imposing goofy transformations. Of course, in the pure SM this conclusion seems somewhat trivial, as $\mu_{H}$ is the only dimensionful parameter. The true power of goofy transformations is that they will also dictate the possible (potentially finite) corrections to $\mu_{H}$ in case other mass scales are present, i.e. help in addressing the “extrinsic” hierarchy problem of the SM [10]. From a simple diagrammatic argument, we expect that finite contributions to $\mu_{H}$ can arise only if the corresponding scale can, in principle, enter the beta function of $\mu_{H}$. If the goofy transformation is imposed, the RG corrections can vanish, which does not automatically preclude finite contributions. Finite contributions to the Higgs mass can be computed from the full effective potential. The Coleman-Weinberg dynamical generation of scale here is also controlled by goofy transformations, which might be especially interesting taking into account their flavor non-universal variants.
@@
We will not discuss UV completions here, but use a spurion argument about new scales to show how goofy constraints play a role. For example, a potentially high scale generated by a condensing scalar field would transform as a spurion depending on the goofy transformation behavior of that scalar. Regarding the Higgs mass term, naive dimensional power counting suggests that any scale should enter the beta functions. However, whether or not it actually does, depends on whether its contribution can be supported by corresponding covariant transforming combination of other couplings (see [3]). If the spurion does not enter the infinite parts contributing to RG running, i.e. the beta function, as one can understand diagrammatically, it can be argued that it should also not enter as a finite contribution.
The more common case is the following. The spurion scale does generally enter the beta function, but its contribution to the beta function vanishes once an appropriate goofy transformation is imposed. In this case, the spurionic scale might still enter in finite contributions. This is the phenomenologically more interesting case, as the electroweak scale needs to be generated after all.
A scale that talks to the Higgs mass unsuppressedly, would have to come from goofy-breaking dynamics ending up in a goofy covariant spurion. If the generated scale does not directly break the goofy transformation, goofy symmetric boundary conditions are possible and natural, and finite corrections to $\mu_{H}^{2}$ , if possible in the first place, only arise suppressed by correct combinations of covariantly transforming softly goofy breaking couplings, namely the kinetic terms and, depending on model building choices, also specific combination of gauge and Yukawa couplings.
@@
In absence of exotic space-time symmetries, one has to accept the fact that goofy transformations are necessarily explicitly (possibly softly) broken by the presence of (flavor universal, canonical) kinetic terms. The goofy transformation is never imposed as an invariance of the full effective action, meaning also that the transformation behavior of a cut-off or renormalization scale is not dictated. Hence, scale generating spurions of various transformation behavior may be considered.
@@
In the SM EFT this demands for an additional power counting of scales, depending on their covariant goofy transformation behavior. In some sense this is analogous to helicity counting [20] for understanding the presence of “holomorphic zeros” in the (non-)renormalization of certain operators [21], [^10] While this has not been shown, we strongly suspect that also helicity counting can be formulated based on an outer automorphism argument, just like for the case of goofy transformations [3, 6].. For example, scales generating effective operators of the type $|H|^{2+4n}$ require a violation of goofy symmetry and will give a boundary condition to feed into the Higgs mass at tree level, while scales that generate operators of the type $|H|^{4n}$ would – if they actually enter the Higgs mass, which depends on details of the model – give a boundary condition that is suppressed by the higher order, potentially sparse (since necessarily goofy breaking), radiative corrections  [^11] As was pointed out to me by Stefan Antusch, also constraints from goofy transformation regarding dimension-0 operators, i.e. vacuum energy, may be considered with potentially interesting conclusions about the cosmological constant problem, and this shall be explored in future work..
Lastly, we note that for non-trivially transforming fermion kinetic terms, the goofy transformation of the gauge coupling terms can be rolled over into a relative sign flip of certain gauge couplings. In two of our examples, this is the case exclusively for the electromagnetic and strong gauge couplings of the RH quarks. For flavor non-universal cases also inter-generational relative sign flips are possible. Currently we are not aware of any observable that is sensitive to either of such relative gauge coupling sign flips, but we think it would be very interesting to find some. Radiative corrections are organized according to such sign flips, irrespectively of whether or not the goofy transformation is imposed.
@@ Acknowledgments
This work is supported by the Portuguese Fundação para a Ciência e a Tecnologia (FCT) through project 2023.06787.CEECIND and contract 2024.01362.CERN, partially funded through POCTI (FEDER), COMPETE, QREN, PRR, and the EU.
Note added.– While I was openly discussing the idea of goofy transformations for the hierarchy problem and consequences, Ref. [24] appeared, which to some extent forced me to release my findings at the present stage. Where overlapping, our conclusions agree.

### v2

@@ Abstract
All of these solutions require a breaking of the protecting symmetry in order to be phenomenologically viable in the sense that they need to allow for electroweak symmetry breaking and the observed Higgs mass. The mechanism of breaking typically renders the Higgs mass computable and introduces deviations in Higgs phenomenology as compared to the plain SM. Also in goofy situations phenomenological consequences are computable in specific models and will lead to testable predictions at future colliders and Higgs factories.
@@
As a model building choice, one possibility to compensate the unusual transformation behavior of gauge-kinetic terms under goofy transformations is by introducing an equally goofy transformation of space-time and gauge fields [1],
$\partial_{\mu}\,\mapsto\,-\mathrm{i}\partial_{\mu}\;,$ and $A_{\mu}\,\mapsto\,-\mathrm{i}A_{\mu}\;.$ (1)
The combined transformation then acts as a regular symmetry, but only leaves invariant the full effective potential if one simultaneously performs a sign flip of the squared UV cutoff, or squared $\overline{\mathrm{MS}}$ renormalization scale [7, 4] (see also [8, 9]). This might be important to take into account for the “intrinsic” hierarchy problem of the SM [10], as a spurious transformation behavior of the cutoff under goofy transformations can determine whether or not, and at what order, it contributes to the otherwise prohibited Higgs quadratic term.
An alternative route to understand the all-order RG stability of the goofy parameter relations – without imposing the exotic space-time transformations in (1) – stems from the observation that the breaking of goofy symmetry by the gauge-kinetic terms can be soft, in the sense that it does not enter quantum corrections of other goofy non-invariant operators, which also explains their RG stability to all orders [3]. Whether or not the explicit breaking by the gauge-kinetic terms actually is soft for a specific goofy transformation, so far has been determined on a case by case basis using explicit higher-loop computations or the all-order arguments of [3, 6]. Some generally applicable arguments are given in footnote  [^1] The gauge-kinetic terms are dimension-four, naively marginal operators, hence, might be expected to lead to a hard symmetry breaking. However, it has been shown in [3] that this is not necessarily the case. Specifically, it was empirically found in the 2HDM that cases where all gauge-kinetic terms break the goofy transformation (“global-sign-flipping goofy transformations”) lead to radiatively stable goofy parameter conditions (i.e. the global sing-flipping gauge-kinetic terms are only softly breaking the goofy symmetry), while “relatively-sign-flipping” goofy transformations of the gauge-kinetic terms lead to a hard breaking of the goofy transformations (i.e. the goofy parameter conditions are then not radiatively stable). On the one hand, these statements straightforwardly generalize to other models also including fermion gauge-kinetic terms, at least if the new particles share a gauge interactions with the scalars. On the other hand, the statements about radiative stability for global-sign-flipping transformations break down if there appear goofy-invariant trilinear couplings (such as, for example, a trilinear scalar coupling). A generally applicable theorem for arbitrary models seems possible but has not been formulated.. Unlike in the case of additional space-time symmetries, in this picture, nothing has to be imposed about the transformation behavior of the RG scale or cutoff, and physical scales can exist which do not have the correct spurion transformation behavior to enter the Higgs quadratic term, hence, giving way for a stabilization of a hierarchy of scales.
@@
[…]
for flavor- or CP-type transformations, respectively  [^3] It is left implicit throughout this work, but for CP-type transformation we always additionally perform the transformation $(t,x)\mapsto(t,-x)$ as well as the usual action of the complex conjugation outer automorphism on gauge fields and fermions [26, 27]. For fermions the CP transformation on flavor and spinor indices reads, for example, $Q_{\mathrm{L}}\mapsto C_{Q_{\mathrm{L}}}\,\mathcal{C}\,Q_{\mathrm{L}}^{*}$, where $\mathcal{C}$ is the charge conjugation matrix and $C_{Q_{\mathrm{L}}}$ a possible flavor-space transformation.. The phases $\alpha$ and $\beta$ here can always be modified by performing an additional global hypercharge transformation and we often adopt a convention such as to set $\alpha=\beta=0$ for the Higgs transformation  [^4] We remark that an independent passive field redefinition of $H$ and $\widetilde{H}$ cannot be used to absorb the phases in Eq. (2) since such an operation commutes with this transformation. Passive field redefinitions, however, can be used to absorb the phases of the CP-type transformation in Eq. (3), which can be used to show that it is basis-change equivalent to an order $2$ transformation. Moreover, passive “goofy” field redefinitions can be used to change to a basis with different sign for the kinetic term..
@@
Extending goofy transformations to fermions, there is a qualification necessary, depending on if one requests the additional space-time symmetries (1) or not. For the scalar kinetic term, the requirement of hermiticity in the case of no additional space-time transformation imposes the very same requirement as demanding invariance in the presence of the additional space-time transformations, Eq. (1). By contrast, for the fermion kinetic term, these two requirements do not align. For example, consider a generic fermion kinetic term (eventually we will take $\Psi=Q_{\mathrm{L}},\,u_{\mathrm{R}},\,d_{\mathrm{R}}$)
@@
where we have included $3\times 3$ unitary transformation matrices in flavor space $A_{\Psi}$, $B_{\Psi}$, $C_{\Psi}$, and $D_{\Psi}$  [^7] See [3] for details in the analogous case of scalars and keep in mind footnote [13]..
Starting from the canonical basis, and without additional exotic space-time transformations, hermiticity of the kinetic term requires $B_{\Psi}^{\dagger}A_{\Psi}=\pm\mathbbm{1}$ or $D_{\Psi}^{\dagger}C_{\Psi}=\pm\mathbbm{1}$ for flavor universal regular $(+)$ or goofy ()(-) transformations, respectively. Different transformations for different $\Psi$’s, and also flavor non-universal transformations are possible. By contrast, superimposing the additional exotic space-time symmetry (1), the overall invariance of the kinetic term requires $B_{\Psi}^{\dagger}A_{\Psi}=\mathrm{i}\mathbbm{1}$ or $D_{\Psi}^{\dagger}C_{\Psi}=\mathrm{i}\mathbbm{1}$, which in this case has to hold universally for all $\Psi=Q_{\mathrm{L}},\,u_{\mathrm{R}},\,d_{\mathrm{R}}$. Obviously, these are very different constraints that can lead to different possibilities regarding the possibly allowed transformations and the resulting phenomenology.
For the case of no extra space-time symmetry there is a lot of model building freedom. In particular, which of the fermion kinetic terms transforms non-trivially and, therefore, explicitly breaks a given goofy transformation, depends on the precise charge assignments under the transformation.
@@
can ensure invariant Yukawa couplings, while the Higgs mass term is prohibited. The transformation is explicitly broken by the $H$, $u_{\mathrm{R}}$ and $d_{\mathrm{R}}$ kinetic terms. The explicit breaking by kinetic terms is not automatically bad, since such a breaking can be soft, see [3] and footnote [11]. Irrespective of whether the breaking is hard or soft, the goofy transformation and its breaking control radiative corrections in potentially phenomenologically interesting ways [3].
@@
Let us also consider the case of imposing the additional space-time transformation, Eq. (1). If this is combined with the transformations (2) or (3) and demanding invariance of the fermion kinetic terms, the flavor universal fermion goofy transformations are limited to
@@
Discussion.– All of our examples show how goofy transformations can be invoked to prohibit the Higgs quadratic term while permitting SM Yukawa couplings. Even if a goofy transformation is explicitly broken by kinetic terms, it can give rise to all-order RG stable fixed points, including mass terms (see footnote [11]). This has been explicitly shown for some cases of the 2HDM in [1, 3], where RG stable fixed points exist even if the canonical gauge-kinetic terms explicitly (softly) break the corresponding goofy transformation  [^9] We add to [3] that for all cases with exact protection of mass terms, one can always find a full rank flavor non-universal (hence non-canonical) kinetic term (this specifies a different theory!) which is actually conserved under a (in that case Higgs flavor dependent) goofy transformation. This implies that the theory with invariant but non-canonical kinetic term, is actually a non-goofy theory - as the kinetic term is invariant - just written in a non-canonical basis. Clearly a (non-)invariant transformation behavior of any term in an irrep is a basis invariant feature.. Irrespective of whether or not a goofy transformation is actually imposed, the transformation acts as an outer automorphism and, therefore, provides a sufficient condition for a symmetry of the system of beta functions that can be larger than the symmetries of the effective action [3, 6].
@@
This implies that a vanishing beta function of $\mu_{H}$ can be stabilized to all order by imposing a goofy transformation of the SM Higgs field. Of course, in the pure SM this conclusion seems somewhat trivial, as $\mu_{H}$ is the only dimensionful parameter. The true power of goofy transformations is that they will also dictate the possible (potentially finite) corrections to $\mu_{H}$ in case other mass scales are present in extensions of the SM or taking into account gravity, i.e. help in addressing the “extrinsic” hierarchy problem of the SM [10].
While goofy transformations are instrumental to prohibit the Higgs quadratic term to begin with, we stress that they do not seem to dictate a unique route along which the Higgs mass and EW scale is re-introduced. In this sense, the possible model building landscape may be as vast as, for example, in supersymmetric solutions to the hierarchy problem. Hence, it is not possible to explore the full landscape of possibilities in the course of just this letter. Rather, we would like to close by giving a few universal arguments that will hold true irrespective of the concrete scenario.
From a simple diagrammatic argument, we expect that finite contributions to $\mu_{H}$ can arise only if the corresponding scale can, in principle (i.e. before imposing goofy parameter relations), enter the beta function of $\mu_{H}$. Once a goofy transformation is imposed to prohibit $\mu_{H}$ and its associated counter-term, the RG corrections must vanish, which does not automatically preclude finite contributions to vanish (finite contributions have been explicitly computed in [9]). Finite contributions to the Higgs mass can be computed from the full effective potential. The Coleman-Weinberg dynamical generation of scale here is also controlled by goofy transformations, which can be especially interesting taking into account their flavor non-universal variants.
@@
We will not discuss UV completions here, but use a spurion argument about new scales to show how goofy constraints play a role. Regarding the Higgs mass term, naive dimensional power counting suggests that any scale should enter the beta functions. However, whether or not it actually does, depends on whether its contribution can be supported by corresponding covariant transforming combination of other couplings (see [3]). For example, a new high scale that transforms invariant under goofy transformations can affect the Higgs mass term only if an appropriate combination of other couplings – necessarily goofy-odd in combination – allows to support the contribution to the beta function. This implies that once the goofy symmetry is imposed, such a combination of couplings is absent and the Higgs mass insensitive to the new high scale. However, things are more involved, as we would like to keep the kinetic terms intact, which are necessarily explicitly goofy-breaking at least for the Higgs field. As we stressed already above, this does not automatically mean the sensitivity to the high scale is restored, as soft symmetry breaking by kinetic terms is a possibility that has been proven to exist.
Another possibility is that we have a high scale spurion that transforms odd under the goofy transformation. This can be the case, for example, if a new scalar field that itself transforms goofy-odd condenses to spontaneously break the goofy transformation. Such a spontaneous breaking can be triggered by a goofy-even mass scale, or via dimensional transmutation á la Coleman-Weinberg. Such a spurion could couple to the Higgs mass term potentially already at the tree level, as the portal term of the new scalar to the Higgs would be allowed by the goofy transformation. This would render the Higgs mass computable (i.e. still prohibit the Higgs mass counter-term), but suggests that the scale of spontaneous goofy breaking should not be too far away from the electroweak scale, for as not to generate a new hierarchy problem in itself.
We stress that goofy symmetry must be broken in some way, as the electroweak scale needs to be generated after all. Again, this bears some similarity to supersymmetry, which also has to be broken in some sequestered sector, and the breaking subsequently mediated to the Higgs to generate a viable EW scale. This supports the conclusion that also in the case of goofy protection of the EW scale, there is most likely a rich zoo of models for goofy symmetry breaking and mediation. This explains why we only briefly mention the spurion analysis, similar to soft supersymmetry breaking.
@@
In absence of exotic space-time symmetries, one has to accept the fact that goofy transformations are necessarily explicitly (possibly softly) broken by the presence of (flavor universal, canonical) kinetic terms. The goofy transformation is never imposed as an invariance of the full effective action, meaning also that the transformation behavior of a cut-off or renormalization scale is not dictated. Hence, scale generating spurions of various transformation behavior can be considered.
@@
In the SM EFT this demands for an additional power counting of scales, depending on their covariant goofy transformation behavior. In some sense this is analogous to helicity counting [21] for understanding the presence of “holomorphic zeros” in the (non-)renormalization of certain operators [22], [^11] While this has not been shown, we strongly suspect that also helicity counting can be formulated based on an outer automorphism argument, just like for the case of goofy transformations [3, 6].. For example, scales generating effective operators of the type $|H|^{2+4n}$ require a violation of goofy symmetry and will give a boundary condition to feed into the Higgs mass at tree level, while scales that generate operators of the type $|H|^{4n}$ would – if they actually enter the Higgs mass, which depends on details of the model – give a boundary condition that is suppressed by the higher order, potentially sparse (since necessarily goofy breaking), radiative corrections  [^12] As was pointed out to me by Stefan Antusch, also constraints from goofy transformation regarding expectation values of dimension-$4$ operators (dimension 0 in the couplings), i.e. vacuum energy, may be considered with potentially interesting conclusions about the cosmological constant problem, and this shall be explored in future work..
Lastly, we note that for non-trivially transforming fermion kinetic terms, the goofy transformation of the gauge coupling terms can be rolled over into a relative sign flip of certain gauge couplings. In two of our examples, this is the case exclusively for the electromagnetic and strong gauge couplings of the RH quarks. For flavor non-universal cases also inter-generational relative sign flips are possible. Currently we are not aware of any observable that is sensitive to either of such relative gauge coupling sign flips, but we think it would be very interesting to find some. Radiative corrections are organized according to such goofy sign flips (and outer automorphisms in general), irrespective of whether or not the according transformation is imposed as a symmetry.
@@ Acknowledgments
This work is supported by the Portuguese Fundação para a Ciência e a Tecnologia (FCT) through projects 2023.06787.CEECIND, UID/00777/2025, and contract 2024.01362.CERN, partially funded through POCTI (FEDER), COMPETE, QREN, PRR, and the EU.
Note added.– While I was openly discussing the idea of goofy transformations for the hierarchy problem and consequences, Ref. [25] appeared, which to some extent forced me to release my findings at the present stage. While ref. [25] focuses on one case of a non-CP-type goofy transformation and imposes the presence of the additional space-time transformation, Eq. (1), I also consider cases with no additional space-time transformations and CP-type transformations, and show that they necessarily have to have some level of flavor dependence. Where overlapping, our conclusions agree. Ref. [25], furthermore, provides some simple explicit extensions of the (goofy symmetric) SM that allows them to reintroduce the Higgs quadratic term and EW symmetry breaking. These are good examples for goofy-symmetric scenarios to address the EW hierarchy problem obeying the general arguments provided in my discussion section.